Exercise 5-5 (Algo) Product and Customer Profitability Analysis [LO5-4, LO5-5]

Thermal Rising, Inc., makes paragliders for sale through specialty sporting goods stores. The company has a standard paraglider model, but also makes custom-designed paragliders. Management has designed an activity-based costing system with the following activity cost pools and activity rates:

Activity Cost Pool Activity Rate
Supporting direct labor $ 16 per direct labor-hour
Order processing $ 184 per order
Custom design processing $ 256 per custom design
Customer service $ 430 per customer

Management would like an analysis of the profitability of a particular customer, Big Sky Outfitters, which has ordered the following products over the last 12 months:

Standard
Model
Custom
Design
Number of gliders 13 3
Number of orders 2 3
Number of custom designs 0 3
Direct labor-hours per glider 28.50 32.00
Selling price per glider $ 1,625 $ 2,360
Direct materials cost per glider $ 476 $ 576

The company’s direct labor rate is $20 per hour.

Required:

Using the company’s activity-based costing system, compute the customer margin of Big Sky Outfitters. (Round your intermediate calculations and final answer to the nearest whole dollar amount. Loss amounts should be entered with a minus sign.)

Customer Margin:
Answer #1
Revenue 28205 =(1625*13)+(2360*3)
Costs:
Direct materials 7916 =(476*13)+(576*3)
Direct labor 9330 =(28.5*20*13)+(32*20*3)
Supporting Direct labor 7464 =(28.5*16*13)+(32*16*3)
Order processing 920 =(2*184)+(3*184)
Custom design processing 768 =(3*256)
Customer service 430
Total costs 26828
Customer margin 1377